# Stationary solutions to stochastic 3D Euler equations in Hölder space

### v1

@@ Appendix 1 A Beltrami waves
Appendix 2 B Estimates of $a_{(\xi)}$ in the proof of Proposition 3.2
For completeness, we include here the detailed proof of the estimates (4.14) and (4.15). We first aim at estimating the $C^{N}_{t,x}$-norm of $\rho=\sqrt{|\mathring{R}_{\ell}|^{2}+\ell^{2}}+c_{*}\zeta_{\ell}$ for $N\in\mathbb{N}$. To this end, we first apply the chain rule [BDLIS15, Proposition C.1] to the function $\Psi(z)=\sqrt{\ell^{2}+z^{2}}$, since $|D^{m}\Psi(z)|\lesssim\ell^{-m+1}$, for $m\geqslant 1$, we obtain for any $N\in\mathbb{N}$
$\|\rho\|_{C^{N}_{t,x}}$ $\lesssim\left\|\sqrt{\ell^{2}+|\mathring{R}_{\ell}|^{2}}\right\|_{C^{0}_{t,x}}+\|D\Psi\|_{C^{0}}\|\mathring{R}_{\ell}\|_{C^{N}_{t,x}}+\|D\Psi\|_{C^{N-1}}\|\mathring{R}_{\ell}\|^{N}_{C^{1}_{t,x}}+\|\zeta_{\ell}\|_{C^{N}_{t}}$ (B.1)
$\lesssim\ell+\|\mathring{R}_{q}\|_{C^{0}_{[t-1,t+1],x}}+\ell^{-N}\|\mathring{R}_{q}\|_{C^{0}_{[t-1,t+1],x}}+\ell^{-2N+1}\|\mathring{R}_{q}\|_{C^{0}_{[t-1,t+1],x}}^{N}+\ell^{-N}\delta_{q+1}\bar{e},$
where we have used definition of $\zeta_{\ell}$ (4.10) in the last inequality. Let us now estimate the $C^{N}_{t,x}$-norm of $a_{(\xi)}$. By Leibniz rule, we get
$\|a_{(\xi)}\|_{C^{N}_{t,x}}\lesssim\sum_{m=0}^{N}\|\rho^{\frac{1}{2}}\|_{C^{m}_{t,x}}\left\|\gamma_{\xi}\left(\textrm{Id}-\frac{\mathring{R}_{\ell}}{\rho}\right)\right\|_{C^{N-m}_{t,x}},$ (B.2)
and estimate each term separately. First, by (4.10)
$\|\rho^{\frac{1}{2}}\|_{C^{0}_{t,x}}\lesssim\|\mathring{R}_{q}\|_{C^{0}_{[t-1,t+1],x}}^{\frac{1}{2}}+\ell^{\frac{1}{2}}+\delta_{q+1}^{\frac{1}{2}}\bar{e}^{\frac{1}{2}},$
and by Lemma A.2
$\left\|\gamma_{\xi}\left(\textrm{Id}-\frac{\mathring{R}_{\ell}}{\rho}\right)\right\|_{C^{0}_{t,x}}\lesssim 1.$
Second, applying [BDLIS15, Proposition C.1] to the function $\Psi(z)=z^{1/2},$ $|D^{m}\Psi(z)|\lesssim|z|^{1/2-m},$ for $m=1,\dots,N$, and using (B.1) and $\rho\geqslant\ell$ and $\ell^{-1}\geqslant\bar{e}$ we obtain for $m\geqslant 1$
$\|\rho^{1/2}\|_{C^{m}_{t,x}}\lesssim$ $\|\rho^{1/2}\|_{C^{0}_{t,x}}+\ell^{-1/2}\|\rho\|_{C^{m}_{t,x}}+\ell^{1/2-m}\|\rho\|_{C^{1}_{t,x}}^{m}$ (B.3)
$\lesssim$ $\|\mathring{R}_{q}\|_{C^{0}_{[t-1,t+1],x}}^{\frac{1}{2}}+\ell^{\frac{1}{2}}+\delta_{q+1}^{\frac{1}{2}}\bar{e}^{\frac{1}{2}}+\ell^{\frac{1}{2}-2m}\|\mathring{R}_{q}\|^{m}_{C^{0}_{[t-1,t+1],x}}$
$+\ell^{-\frac{1}{2}-m}\|\mathring{R}_{q}\|_{C^{0}_{[t-1,t+1],x}}+\ell^{\frac{1}{2}-2m}\delta_{q+1}^{m}+\ell^{-\frac{1}{2}-m}\delta_{q+1}.$
For $m\geqslant 1$, using $\delta_{q+1}\leqslant 3rL^{2}$ the above is bounded by
$\ell^{\frac{1}{2}-2m}(3rL^{2}+\|\mathring{R}_{q}\|_{C^{0}_{[t-1,t+1],x}})^{m}.$
We proceed with a bound for $\left\|\gamma_{\xi}\left(\textrm{Id}-\frac{\mathring{R}_{\ell}}{\rho}\right)\right\|_{C^{N-m}_{t,x}}$ for $m=0,\dots,N-1$. Keeping [BDLIS15, Proposition C.1] in mind, we need to estimate
$\left\|\frac{\mathring{R}_{\ell}}{\rho}\right\|_{C^{N-m}_{t,x}}+\left\|\frac{\nabla_{t,x}\mathring{R}_{\ell}}{\rho}\right\|_{C^{0}_{t,x}}^{N-m}+\left\|\frac{\mathring{R}_{\ell}}{\rho^{2}}\right\|_{C^{0}_{t,x}}^{N-m}\|\rho\|_{C^{1}_{t,x}}^{N-m}.$ (B.4)
We use $\rho\geqslant\ell$ to have
$\left\|\frac{\nabla_{t,x}\mathring{R}_{\ell}}{\rho}\right\|_{C^{0}_{t,x}}^{N-m}\lesssim\ell^{-2(N-m)}\|\mathring{R}_{q}\|^{N-m}_{C_{[t-1,t+1],x}^{0}},$ (B.5)
and in view of $|\frac{\mathring{R}_{\ell}}{\rho}|\leqslant 1$
$\left\|\frac{\mathring{R}_{\ell}}{\rho^{2}}\right\|_{C^{0}_{t,x}}^{N-m}\lesssim\left\|\frac{1}{\rho}\right\|_{C^{0}_{t,x}}^{N-m}\lesssim\ell^{-(N-m)},$
and by (B.1) and $\delta_{q+1}\leqslant 3rL^{2}$
$\|\rho\|_{C^{1}_{t,x}}^{N-m}\lesssim\ell^{-(N-m)}\|\mathring{R}_{q}\|^{N-m}_{C_{[t-1,t+1],x}^{0}}+\ell^{-(N-m)},$
hence
$\left\|\frac{\mathring{R}_{\ell}}{\rho^{2}}\right\|_{C^{0}_{t,x}}^{N-m}\|\rho\|_{C^{1}_{t,x}}^{N-m}\lesssim\ell^{-2(N-m)}\|\mathring{R}_{q}\|^{N-m}_{C_{[t-1,t+1],x}^{0}}+\ell^{-2(N-m)}.$ (B.6)
To estimate the first term in (B.4), we write
$\left\|\frac{\mathring{R}_{\ell}}{\rho}\right\|_{C^{N-m}_{t,x}}\lesssim\sum_{k=0}^{N-m}\|\mathring{R}_{\ell}\|_{C^{k}_{t,x}}\left\|\frac{1}{\rho}\right\|_{C^{N-m-k}_{t,x}},$ (B.7)
where for $N-m-k=0$ we have
$\left\|\frac{1}{\rho}\right\|_{C^{0}_{t,x}}\lesssim\ell^{-1},$
and for $k=0,\dots,N-m-1$ using (B.1)
$\left\|\frac{1}{\rho}\right\|_{C^{N-m-k}_{t,x}}$ $\lesssim\left\|\frac{1}{\rho}\right\|_{C^{0}_{t,x}}+\ell^{-2}\|\rho\|_{C^{N-m-k}_{t,x}}+\ell^{-(N-m-k){-1}}\|\rho\|_{C^{1}_{t,x}}^{N-m-k}$
$\lesssim\ell^{-2(N-m-k)-1}+\ell^{-(N-m-k)-2}\|\mathring{R}_{q}\|_{C_{[t-1,t+1],x}^{0}}+\ell^{-2(N-m-k)-1}\|\mathring{R}_{q}\|_{C_{[t-1,t+1],x}^{0}}^{N-m-k}.$
Altogether, we therefore bound (B.7) as
$\left\|\frac{\mathring{R}_{\ell}}{\rho}\right\|_{C^{N-m}_{t,x}}\lesssim\ell^{-2(N-m)-1}$ $\|\mathring{R}_{q}\|_{C_{[t-1,t+1],x}^{0}}+\ell^{-(N-m)-2}\|\mathring{R}_{q}\|^{2}_{C_{[t-1,t+1],x}^{0}}$ (B.8)
$+\ell^{-2(N-m)-1}\|\mathring{R}_{q}\|^{N-m+1}_{C_{[t-1,t+1],x}^{0}}.$
Finally, plugging (B.5), (B.6) and (B.8) into (B.4) leads to
$\left\|\gamma_{\xi}\left(\textrm{Id}-\frac{\mathring{R}_{\ell}}{\rho}\right)\right\|_{C^{N-m}_{t,x}}$ $\lesssim\ell^{-2(N-m)-1}\|\mathring{R}_{q}\|_{C_{[t-1,t+1],x}^{0}}+\ell^{-(N-m)-2}\|\mathring{R}_{q}\|^{2}_{C_{[t-1,t+1],x}^{0}}$
$+\ell^{-2(N-m)}\|\mathring{R}_{q}\|^{N-m}_{C_{[t-1,t+1],x}^{0}}+\ell^{-2(N-m)-1}\|\mathring{R}_{q}\|^{N-m+1}_{C_{[t-1,t+1],x}^{0}}+\ell^{-2(N-m)}.$
For $N-m\geqslant 1$ the above is bounded by
$\ell^{-2(N-m)-1}(\|\mathring{R}_{q}\|^{N-m+1}_{C_{[t-1,t+1],x}^{0}}+\|\mathring{R}_{q}\|_{C_{[t-1,t+1],x}^{0}}+1).$
@@
Combining this with the bounds (B.3) above and plugging into (B.2) yields for $N\geqslant 2$
@@
where we used inductive assumption (3.8d) for $\mathring{R}_{q}$ in the last inequality. Moreover, for $N=1$, by (B.3) and chain rule, we obtain
$\|a_{(\xi)}\|_{C^{1}_{t,x}}$ $\lesssim\|\rho^{\frac{1}{2}}\|_{C^{1}_{t,x}}+\|\rho^{\frac{1}{2}}\|_{C^{0}_{t,x}}\left\|\gamma_{\xi}\left(\mathrm{Id}-\frac{\mathring{R}_{\ell}}{\rho}\right)\right\|_{C^{1}_{t,x}}$ (B.10)
$\lesssim\ell^{-\frac{3}{2}}(\delta_{q+1}+\|\mathring{R}_{q}\|_{C_{[t-1,t+1],x}^{0}})+(\delta_{q+1}^{\frac{1}{2}}+\ell^{\frac{1}{2}}+\|\mathring{R}_{q}\|_{C_{[t-1,t+1],x}^{0}}^{\frac{1}{2}})\left\|\frac{\mathring{R}_{\ell}}{\rho}\right\|_{C^{1}_{t,x}}.$
Since $\rho\geqslant\ell$ and $|\frac{\mathring{R}_{\ell}}{\rho}|\leqslant 1$, by (B.3) and using chain rule again, we obtain
$\left|\nabla_{t,x}\left(\frac{\mathring{R}_{\ell}}{\rho}\right)\right|\lesssim\frac{|\nabla_{t,x}\mathring{R}_{\ell}|}{\rho}+\frac{|\mathring{R}_{\ell}\cdot\nabla_{t,x}\rho|}{\rho^{2}}\lesssim\ell^{-2}\|\mathring{R}_{q}\|_{C^{0}_{[t-1,t+1],x}}+\ell^{-2}.$ (B.11)
Pluging (B.11) into (B.10), we obtain
$\|a_{(\xi)}\|_{C^{1}_{t,x}}\lesssim\ell^{-2}(1+\|\mathring{R}_{q}\|_{C^{0}_{[t-1,t+1],x}})^{\frac{3}{2}}\leqslant\ell^{-2}\lambda_{q+3}^{2\gamma}.$ (B.12)
And by the definition of $a_{(\xi)}$, it is easy to get
@@
Appendix 3 C Estimates for transport equantions
@@
We recall the following results from [BDLIS15, Proposition D.1, (133), (134)]: {equ} ∥f(t)∥_C_x^1⩽∥f_0∥_C_x^1e^(t-t_0)∥v∥_C^0_[t_0,t]C^1_x+∫_t_0^te^(t-τ)∥v∥_C_[t_0,t]^0C_x^1∥g(τ)∥_C_x^1 dτ, and more generally, for any $N\geqslant 2$, there exists a constant $C=C_{N}$ such that
@@
We aim at estimating the $C_{x}^{0}$-norm of $\nabla\Phi(t)-\textrm{Id}$. First, let $\Psi(s,x)=\Phi(s,x)-x$, then $\Psi$ statifies the following equation:
@@
hence (4.8a) holds ture. Then, (4.8b) follows from (C.5), we have
@@
Similarly, we use (C.2) and (C.4) again, we obtain for any $n\geqslant 1$
@@
Differentiating both sides of (4.7) and using (3.8a), (3.8c), (4.31), (C.6) and (C.7), we can also obtain

### v3

@@ Appendix 1 A Beltrami waves
Appendix 2 B Proof of Proposition 5.1
We concisely present the proof of Proposition 5.1, following closely the calculation of [HZZ22b, Appendix B]. We only write the necessary changes compared to [HZZ22b, Appendix B]. By the chain rule, we have
$\|a_{(\xi)}\|_{C^{N}_{t,x}}\lesssim\sum_{m=0}^{N}\|\rho^{\frac{1}{2}}\|_{C^{m}_{t,x}}\big{\|}\gamma_{\xi}\big{(}\textrm{Id}-\rho^{-1}\mathring{R}_{\ell}\big{)}\big{\|}_{C^{N-m}_{t,x}}.$ (B.1)
We estimate the two terms on the right hand side in (B.1) separately. First, using Lemma A.2 and (4.6) we have
$\|\rho^{\frac{1}{2}}\|_{C^{0}_{t,x}}\lesssim\|\mathring{R}_{q}\|_{C^{0}_{[t-1,t+1],x}}^{\frac{1}{2}}+\ell^{\frac{1}{2}}+\delta_{q+1}^{\frac{1}{2}}\bar{e}^{\frac{1}{2}},\qquad\|\gamma_{\xi}(\textrm{Id}-\rho^{-1}\mathring{R}_{\ell})\|_{C^{0}_{t,x}}\lesssim 1.$
Second, we use [HZZ22b, (B.4)] to derive for $m\geqslant 1$
$\|\rho^{1/2}\|_{C^{m}_{t,x}}\lesssim\|\rho^{1/2}\|_{C^{0}_{t,x}}+\ell^{-1/2}\|\rho\|_{C^{m}_{t,x}}+\ell^{1/2-m}\|\rho\|_{C^{1}_{t,x}}^{m}.$ (B.2)
By the definition of $\rho$ in (4.6), it follows from [HZZ22b, (B.1)] that for any $m\geqslant 1$
$\|\rho\|_{C^{m}_{t,x}}\lesssim\ell+\|\mathring{R}_{q}\|_{C^{0}_{[t-1,t+1],x}}+\ell^{-m}\|\mathring{R}_{q}\|_{C^{0}_{[t-1,t+1],x}}+\ell^{-2m+1}\|\mathring{R}_{q}\|_{C^{0}_{[t-1,t+1],x}}^{m}+\ell^{-m}\delta_{q+1}\bar{e}.$ (B.3)
Hence, combining (B.2) with (B.3), we have for any $m\geqslant 1$
$\|\rho^{1/2}\|_{C^{m}_{t,x}}\lesssim\ell^{\frac{1}{2}-2m}(1+\|\mathring{R}_{q}\|_{C^{0}_{[t-1,t+1],x}})^{m}.$ (B.4)
We now consider the second term on the right hand side of (B.1). By [HZZ22b, (B.5)], we have for $m=0,\dots,N-1$
$\left\|\gamma_{\xi}\left(\textrm{Id}-\rho^{-1}\mathring{R}_{\ell}\right)\right\|_{C^{N-m}_{t,x}}\lesssim\left\|\rho^{-1}\nabla_{t,x}\mathring{R}_{\ell}\right\|_{C^{0}_{t,x}}^{N-m}+\left\|\rho^{-2}\mathring{R}_{\ell}\right\|_{C^{0}_{t,x}}^{N-m}\|\rho\|_{C^{1}_{t,x}}^{N-m}+\left\|\rho^{-1}\mathring{R}_{\ell}\right\|_{C^{N-m}_{t,x}}.$ (B.5)
Using $\rho\geqslant\ell$, $|\rho^{-1}\mathring{R}_{\ell}|\leqslant 1$, (B.3) and [HZZ22b, (B.6)], we deduce that for any $N-m\geqslant 1$
$\left\|\gamma_{\xi}\left(\textrm{Id}-\rho^{-1}\mathring{R}_{\ell}\right)\right\|_{C^{N-m}_{t,x}}\lesssim\ell^{-2(N-m)-1}(\|\mathring{R}_{q}\|_{C_{[t-1,t+1],x}^{0}}+1)^{N-m+1}.$ (B.6)
@@
Then combining (B.6) with the bound (B.4) and plugging into (B.1) yields for $N\geqslant 2$
@@
where we used inductive assumption (3.8d) for $\mathring{R}_{q}$ in the last inequality.
Moreover, by a similar calculation as above, we also have
$\|a_{(\xi)}\|_{C^{1}_{t,x}}$ $\lesssim\ell^{-2}(1+\|\mathring{R}_{q}\|_{C^{0}_{[t-1,t+1],x}})^{\frac{3}{2}}\leqslant\ell^{-2}\lambda_{q+3}^{2\gamma},$ (B.8)
@@
Appendix 3 C Estimates for transport equations
@@
We recall the following results from [BDLIS15, Proposition D.1, (133), (134)]: {equ} ∥f(t)∥_C_x^1⩽∥f_0∥_C_x^1e^(t-t_0)∥v∥_C^0_[t_0,t]C^1_x+∫_t_0^te^(t-τ)∥v∥_C_[t_0,t]^0C_x^1∥g(τ)∥_C_x^1 dτ. More generally, for any $N\geqslant 2$, there exists a constant $C=C_{N}$ such that
@@
We aim to estimate the $C_{x}^{0}$-norm of $\nabla\Phi(t)-\textrm{Id}$. First, let $\Psi(s,x)=\Phi(s,x)-x$, then $\Psi$ satisfies the following equation:
@@
hence (4.4a) holds. Then, (4.4b) follows from (C.5), we have
@@
Similarly, we use (C.2) and (C.4) again to obtain for any $n\geqslant 1$
@@
Differentiating both sides of (4.3) and using (3.8a), (3.8c), (5.13), (C.6) and (C.7), we also obtain